Write the isotope needed, by identifying A, Z and X, to balance the following nuclear fusion reaction:


3 3 1 A1H+ 1H→2 0n+ Z X

Express your answer as an isotope using prescripts.

1
Expert's answer
2021-02-26T06:01:55-0500

"\frac{3}{1}H + \frac{3}{1}H = \frac{4}{2}He + 2\frac{1}{0}n"